Consider a planet of mass (m) revolving around the sun. The time period (T) of revolution of the planet depends on the mass of the sun (M), the radius of the orbit (r) and the gravitational constant (G). Verify Kepler's third law of planetary motion


1
Expert's answer
2021-05-03T10:31:54-0400

"\frac{a^3}{T^2(M+m)}=\frac{G}{4\pi^2}=\text{const}."